# Mind-to-Image: Projecting Visual Mental Imagination of the Brain from fMRI

### v1

@@ 1 Introduction
Figure 1: Overview of our method. Left: classic fMRI-to-Image approach where a subject sees images in an MRI scanner, bold data are recorded then used to train a model that convert brain data to match the seen images. Right: in our approach we let the subject imagine previously seen images to collect a dataset, then train a model to reconstruct previously seen images from visual imagination brain data. At inference time we collect data where the subject completely imagines new images based on instructions and use the trained model of the first step to generate the imagined images.
@@
While using models based on reconstruction of images based on visual stimuli to generate images from imagination is theoretically possible, we believe that we need novel approaches to tackle the issue of visualizing imagination in the brain, particularly in data collection and model development. To this end, we introduce a protocol inspired by the NSD dataset but tailored to capture the nuances of visual imagination. Our approach distinguishes between weak and strong imagination, developing specific protocols for each type of visual imagery. We restrict our analysis on two distinct modalities of images: portraits and landscapes, specifically within the theme of surrealism. This choice is not only methodological but also thematic, aligning with the project’s broader artistic aspirations. The intersection of art and science, particularly through the lens of surrealism and its exploration of automatic processes, provides a rich context for our research.
@@ 2 Related work
The development of a comprehensive data collection protocol is one the main contributions of our study, aimed at capturing the complex neural correlates of visual imagination. Drawing inspiration from the NSD paper (Allen et al., 2022), our protocol is designed to explore two distinct facets of visual imagination: weak imagination (imagination from memory) and strong imagination (pure imagination). This dual approach enables us to delve into the nuanced ways in which the brain navigates between recalling visual information and generating novel visual constructs.
@@ 3.1 Weak imagination: data collection protocol
The protocol for weak imagination aims at recording brain activity corresponding to the recall of previously seen images. In this phase, subjects are first shown an image for 3 seconds, followed by a 1-second rest period. This sequence is repeated with three different images, each selected from our curated set of 600 surrealist portraits and 600 surrealist landscapes (including both real and generated images using Versatile Diffusion (Xu et al., 2023) and Midjourney), amounting to a diverse array of visual stimuli. To engage the brain’s imaginative capacities, the same images are then flashed for 0.1 seconds, prompting the subject to imagine the image for 5 seconds before resting for another second. This process is repeated over a span of 6 hours, encompassing a total of 1200 images. Out of the data collected, 100 data points (comprising brain activity and the corresponding image) are earmarked for validation and evaluation, with the remainder serving as the training set for our fMRI-to-image model. The evaluation focuses on the model’s ability to generate images that not only match the correct modality (portrait or landscape) but also reflect the contents of the original image.
@@ 3.2 Strong imagination: data collection protocol
For strong imagination, the process pivots towards the generation of entirely new visual content based on verbal prompts. Subjects are instructed to imagine a portrait or landscape evoking one of ten primal emotions (such as fear, happiness, love, etc.). This task is designed to stimulate the pure imagination process, encouraging the brain to construct novel visual imagery without the reference of previously seen images. Each imagination session lasts for 6 seconds, followed by a 1-second rest, cycling through combinations of landscape/portrait and emotional prompts. This cycle is repeated 10 times. Unlike the data collected for weak imagination, this dataset is solely used during the model’s inference stage. The evaluation here is similar to the weak imagination evaluation: quantitively assess if the model matches the correct modality (portrait or landscape) and qualitatively assess if the contents of the generated image matches the visual imagery of the subject. To ensure that we recall what the subject has imagined, the subject provides an oral description of the visual contents of the imagined image.
@@ 5.2 Can we reconstruct images from weak imagination / memory?
Figure 3: Results for weak imagination on the validation dataset.
In the exploration of weak imagination, where subjects were prompted to recall previously seen images, our results highlight the capacity of the model to identify to the original class of the images. Specifically, if a subject recalled a portrait, the reconstructed image is almost always a portrait, and similarly for landscapes. This observation suggests a the ability of the model to grasp the category of the imagined content accurately within the BOLD data. To quantitatively assess this capability, we fine-tune a pretrained ResNet50 model (He et al., 2016) to classify images as either portraits or landscapes using the images in our dataset. We apply this classifier to the reconstructed images from the weak imagination tasks of the validation set. We obtain a classification accuracy of 91% in correctly identifying the reconstructed images as belonging to their original categories of either portraits or landscapes. This level of accuracy demonstrates the model’s ability in capturing the category information within the neural activity during weak imagination.
@@ 5.3 Can we reconstruct images from strong imagination / pure imagination?
Figure 4: Results from strong imagination.

### v3

@@ 1 Introduction
Figure 1: Overview of our method. Left: classic fMRI-to-Image approach where a subject sees images in an MRI scanner, bold data are recorded then used to train a model that convert brain data to match the seen images. Right: in our approach we let the subject imagine previously seen images from a dataset of surrealist images (face portraits and nature landscapes) which we create. The brain data associated to the recollection of such images is gathered. It is then used to train a model to reconstruct previously seen images from visual imagination brain data. At inference time we collect brain data where the subject completely imagines new images based on instructions and use the trained model of the first step to generate the imagined images.
@@
While using models based on reconstruction of images based on visual stimuli to generate images from imagination is theoretically possible, we believe that we need novel approaches to tackle the issue of visualizing imagination in the brain, particularly in data collection and model development. To this end, we introduce a protocol inspired by the NSD dataset but tailored to capture the nuances of visual imagination. Our approach distinguishes between weak and strong imagination, developing specific protocols for each type of visual imagery. We restrict our analysis on one theme of images: surrealism, and two distinct modalities of images: portraits and landscapes. This is done by collecting a dataset of surrealist face portraits and nature landscapes, using real and generated images. This choice is not only methodological but also thematic, aligning with the project’s broader artistic aspirations. The intersection of art and science, particularly through the lens of surrealism and its exploration of automatic processes, provides a rich context for our research.
@@ 2 Related work
The development of a comprehensive data collection protocol is one the main contributions of our study, aimed at capturing the complex neural correlates of visual imagination. Drawing inspiration from the NSD paper (Allen et al., 2022), our protocol is designed to explore two distinct facets of visual imagination: weak imagination (imagination from memory) and strong imagination (pure imagination). This dual approach enables us to study how the brain navigates between recalling visual information and generating novel visual constructs.
3.1 Image dataset used: Surrealism
We do not use the images from the COCO dataset (Lin et al., 2014) as it is done in the NSD dataset. Because our research is also associated to an artistic research, we decided to narrow the dataset to one theme (surrealism) and two modes (face portraits and nature landscapes). This choice narrows the visual space of the fMRI-to-Image model, which simplify the task and allows us to have results without having to do 40 hours of scans as in the NSD dataset.
Thus, we created our own dataset of images for our experiments. The images come from three sources: real images and generated images using Versatile Diffusion (Xu et al., 2023) and Midjourney. The dataset comprises 1200 images.
This dataset is used in the weak imagination protocol, where the images are recollected from memory and imagined. Then, the images are used to train the fMRI-to-Image model. Finally, only brain data is given to the fMRI-to-Image model to produce the reconstructions, both in the weak and strong imagination cases.
@@ 3.2 Weak imagination: data collection protocol
The protocol for weak imagination aims at recording brain activity corresponding to the recall of previously seen images. In this phase, subjects are first shown an image for 3 seconds, followed by a 1-second rest period. This sequence is repeated with three different images, each selected from our curated set of 600 surrealist portraits and 600 surrealist landscapes. To engage the brain’s imaginative capacities, the same images are then flashed for 0.1 seconds, prompting the subject to imagine the image for 5 seconds before resting for another second. This process is repeated over a span of 6 hours, encompassing a total of 1200 images. Out of the data collected, 100 data points (comprising brain activity and the corresponding image) are earmarked for validation and evaluation, with the remainder serving as the training set for our fMRI-to-image model. The evaluation focuses on the model’s ability to generate images that not only match the correct modality (face portrait or nature landscape) but also reflect the contents of the original image.
@@ 3.3 Strong imagination: data collection protocol
For strong imagination, the process pivots towards the generation of entirely new visual content based on verbal prompts. Subjects are instructed to imagine a portrait or landscape evoking one of ten primal emotions (such as fear, happiness, love, etc.). This task is designed to stimulate the pure imagination process, encouraging the brain to construct novel visual imagery without the reference of previously seen images. Each imagination session lasts for 6 seconds, followed by a 1-second rest, cycling through combinations of landscape/portrait and emotional prompts. This cycle is repeated 10 times. Unlike the data collected for weak imagination, this dataset is solely used during the model’s inference stage. The evaluation here is similar to the weak imagination evaluation: quantitively assess if the model matches the correct modality (face portrait or nature landscape) and qualitatively assess if the contents of the generated image matches the visual imagery of the subject. To ensure that we recall what the subject has imagined, the subject provides an oral description of the visual contents of the imagined image.
@@ 5.2 Can we reconstruct images from weak imagination / memory?
Figure 3: Results for weak imagination on the validation dataset. Top row: images from the validation set of our surrealism image dataset, which were recollected from memory by the subject in the fMRI scanner. Bottom row: reconstructed images from the fMRI-to-Image model based on the brain data associated to the recollection of the images in the top row.
In the exploration of weak imagination, where subjects were prompted to recall previously seen images, our results highlight the capacity of the model to identify to the original class of the images. Specifically, if a subject recalled a face portrait, the reconstructed image is almost always a face portrait, and similarly for nature landscapes. This observation suggests a the ability of the model to grasp the category of the imagined content accurately within the BOLD data. To quantitatively assess this capability, we fine-tune a pretrained ResNet50 model (He et al., 2016) to classify images as either face portraits or nature landscapes using the images in our dataset. We apply this classifier to the reconstructed images from the weak imagination tasks of the validation set. We obtain a classification accuracy of 91% in correctly identifying the reconstructed images as belonging to their original categories of either portraits or landscapes. This level of accuracy demonstrates the model’s ability in capturing the category information within the neural activity during weak imagination.
@@ 5.3 Can we reconstruct images from strong imagination / pure imagination?
Figure 4: Results from strong imagination. The subject is given an instruction on the fMRI screen, such as ”Imagine a portrait representing optimism”. Then, the subject purely imagines such image, giving an oral description. Then, the associated brain data is fed to the fMRI-to-Image model to produce a reconstruction, which is then evaluated compared to the oral description.